# RV-Syn: Rational and Verifiable Mathematical Reasoning Data Synthesis based on Structured Function Library

### v1

@@ Abstract
The advancement of reasoning capabilities in Large Language Models (LLMs) requires substantial amounts of high-quality reasoning data, particularly in mathematics. Existing data synthesis methods, such as data augmentation from annotated training sets or direct question generation based on relevant knowledge points and documents, have expanded datasets but face challenges in mastering the inner logic of the problem during generation and ensuring the verifiability of the solutions. To address these issues, we propose RV-Syn, a novel Rational and Verifiable mathematical Synthesis approach. RV-Syn constructs a structured mathematical operation function library based on initial seed problems and generates computational graphs as solutions by combining Python-formatted functions from this library. These graphs are then back-translated into complex problems. Based on the constructed computation graph, we achieve solution-guided logic-aware problem generation. Furthermore, the executability of the computational graph ensures the verifiability of the solving process. Experimental results show that RV-Syn surpasses existing synthesis methods, including those involving human-generated problems, achieving greater efficient data scaling. This approach provides a scalable framework for generating high-quality reasoning datasets.
@@
Jiapeng Wang1thanks:   Equal contribution., Jinhao Jiang [^1] footnotemark: 1, Zhiqiang Zhang, Jun Zhou, Wayne Xin Zhao1thanks:   Corresponding author., 1Gaoling School of Artificial Intelligence, Renmin University of China 2Ant Group wangjp1010@ruc.edu.cn, batmanfly@gmail.com
@@ 1 Introduction
To address this scarcity, researchers have explored various synthesizing methods, particularly in the mathematics domain. The mainstream methods involve data augmentation based on existing annotated training sets, such as GSM8K Cobbe et al. (2021) and MATH Hendrycks et al. (2021), ranging from self-evolving instructions Xu et al. (2024); Zeng et al. (2024) and question paraphrasing Yu et al. (2024), to solution augmentation Lu et al. (2024). However, these methods are limited by the available training data, constraining the synthesis diversity Li et al. (2024c, 2023). To enhance diversity, recent approaches enable LLMs to generate a large scale of questions from various mathematics-related sources, including web pages Yue et al. (2024) and knowledge points Tang et al. (2024) from web corpora or textbooks.
@@
However, as shown in Figure 1, these direct problem generation methods encounter challenges in lacking a deep understanding and simulation of the underlying reasoning processes. As a result, the synthesized problems may appear superficially plausible but actually suffer from insufficient reasoning depth and logical inconsistencies (such as internal contradictions), especially when generating complex multi-step reasoning problems. Furthermore, a long-standing challenge in synthetic data generation is the difficulty in supporting the validation of solution correctness during the synthesis process, ultimately compromising training efficiency.
@@
Inspired by this, we propose RV-Syn, a novel method for mathematical problem synthesis that emphasizes enhanced rationale and verifiability. This method comprises three stages. First, we extract solution processes from existing problems by prompting the model to generate code-based solutions (computational graphs) where each function performs a specific mathematical operation. Second, a structured function library in graph format is built, detailing the relationships between these functions. Third, a graph-aware sampling strategy combines these functions to create new computational graphs, which are then executed and back-translated to synthesize new problems with correct answers. By generating questions through solution guidance and utilizing the structured and executable nature of code as a blueprint for problem design, RV-Syn effectively provides the rich semantics and control flows required for complex reasoning. As shown in Figure 1, this enables the synthesis of diverse and sophisticated reasoning structures—often challenging to achieve for direct problem generation—including multi-hop reasoning (sequential execution), iterative loops (for-loops), and modular structures (function calls), while naturally providing ground-truth labels.
We conducted extensive experiments to evaluate the proposed RV-Syn method using various LLMs. The results demonstrate that our method achieves superior average performance across five benchmarks compared to existing approaches, including those involving human-designed problems (e.g., Numina-Math). Notably, RV-Syn outperforms the previous state-of-the-art method while utilizing only half the data, leading to a more efficient scaling curve.
@@ 2 Related Work
Researchers have proposed various approaches to enhance the mathematical reasoning capabilities, including methods applied in training or inference stages. During the training stage, existing methods aim to enhance the LLMs from the aspects of pre-training and post-training. Specifically, some studies Shao et al. (2024); Chen et al. (2024a) involve collecting extensive math-related corpora and enhancing the foundational mathematical capabilities of LLMs through continual pre-training. In contrast, other studies Tong et al. (2024); Huang et al. (2024b); Toshniwal et al. (2024) focus on synthesizing a substantial amount of high-quality math-related instructional data, further refining the problem-solving abilities of LLMs through post-training. During the inference stage, there are two primary prompting approaches, including Chain-of-Thought (CoT) Wei et al. (2022) and Program-of-Thought (PoT) Chen et al. (2023). Specifically, some studies Yao et al. (2023); Shinn et al. (2023) utilize CoT to elicit LLMs’ inherent reasoning ability with a step-by-step reasoning process. Furthermore, other studies Gou et al. (2024); Wang et al. (2024); Chen et al. (2024b); Wen et al. (2024) incorporate PoT to allow LLMs to utilize external computors during reasoning. In this work, we focus on the training stage and try to synthesize instruction data by leveraging the strengths of both the CoT and PoT methods.
@@
Existing research has proposed various synthesizing approaches: Firstly, existing studies focus on synthesizing new mathematical problems based on few-shot problems Li et al. (2024a), math-related documents Yue et al. (2024), key knowledge points Huang et al. (2024a), or self-evolution Luo et al. (2023). Furthermore, with the improvement of LLMs’ capabilities, subsequent studies Zhou et al. (2024); Ding et al. (2024) explore to utilize of open-source LLMs (i.e., DeepSeek-Math and Qwen2-Math) to scale the number of synthesized problems. Despite their effectiveness, direct problem generation suffers from a lack of fine-grained control over the problem’s internal logic, leading to lower-quality problems. Besides, these methods frequently lack mechanisms for validating the correctness of generated answers, which may introduce errors and affect the final performance after instruction tuning. In contrast, our method starts from the internal logic of problem-solving, rather than directly synthesizing problems. This provides more refined logical consistency and improves problem quality, while also enabling effective correctness control.
@@ 3.1 Problem Decomposition
We first represent the solutions of existing mathematical problems with sequences of functions into a standardized format shown in Appendix E leveraging LLMs. Each function represents a specific mathematical skill with defined parameters, forming an executable computational graph that captures the data flow of the problem-solving process (examples shown in Figure 2). To ensure the correctness of the extracted solutions and their associated functions, we filter out erroneous cases by executing the solution code using a Python interpreter and comparing its output with the ground-truth answer provided in the corresponding Chain-of-Thought annotation. This results in a large-scale, high-quality set of computational graphs derived from real-world problems.
@@ 3.2 Graph-based Function Library Construction
After decomposing the problem, we extract functions from the computational graphs. To further enhance the quality of the function library, we employ model-based labeling. We utilize LLMs to validate the correctness of the functions and identify their corresponding mathematical topics, eliminating any functions that contain errors. The topic definitions are provided in Table 6, and the prompts are shown in Appendix C. To describe the relationships between these functions, each function is represented as a node, and they are organized into a graph-based function library.
Specifically, the construction process involves two key steps: structure and semantic-aware function merging, followed by co-occurrence and topic-based node connecting.
3.2.1 Structure and Semantic-aware Function Merging
@@
Specifically, we examine the similarity between two functions by analyzing both structure and semantic aspects. For the structure, we focus on the computational operations without considering specific variable names. For instance, expressions such as “$(a+b)/c$” and “$(m+n)/d$” are considered equivalent. To accomplish this, we extract the function body and parse it using Abstract Syntax Trees (ASTs). For the semantic, we compute hash values for the function’s docstrings. If the ASTs or hash values of two functions are identical, we merge these functions into a single node. We show an example of two similar functions in Table 9 at Appendix D. This approach allows similar functions (e.g., matrix multiplication and matrix projection) to preserve their distinct mathematical meanings while minimizing redundancy in the final function library, offering advantages over direct deduplication.
@@ 3.2.2 Co-occurrence and Topic-based Node Connecting
Specifically, for any two given nodes (i.e.,, two sets of functions), we first check whether there exists at least one pair of functions—one from each set—that co-occurred in the same computational graph. If so, we add a co-occurrence-based edge between these two nodes. If not, we then check whether there is at least one pair of functions sharing the same topic. If this condition is met, we add a topic-consistent edge. If neither condition is satisfied, no edge is added between the nodes.
In this way, we finally obtain an incompletely connected graph that contains multiple types of edges, which depicts the relationships between nodes. This serves as the basis for the subsequent recombination of functions. Some statistics of the graph-based function library are shown in Table 6 and 7 in Appendix A.
@@ 3.3 Problem Synthesis
Based on the constructed graph-format function library, we are able to synthesize complex problems in a solution-guided logic-aware approach. This process consists of two primary steps: first sampling various functions to regenerate the computational graph, then synthesizing the final complex problems.
@@ 3.3.1 Computational Graph Regeneration
To balance the reasonableness and novelty of the computational graphs, we design three sampling strategies: first, selecting nodes connected by co-occurrence-based edges; second, selecting nodes connected by topic-based edges; and third, selecting unconnected nodes. During the sampling process, nodes are randomly selected using these strategies. Since each node represents a set of functions, we further randomly select one function per node. Subsequently, we generate a computational graph based on these selected functions as detailed in Appendix C.
This approach offers several advantages: Firstly, by combining functions, the generated computational graph can comprehensively cover mathematical skills while minimizing redundancy. Secondly, the sampling strategy effectively balances high-frequency combinations of mathematical skills (i.e., connected by edges) with long-tail combinations (i.e., without connected edges). In our setup, we randomly select from the three strategies when synthesizing each data sample.
3.3.2 Problem Back-translation
After obtaining the new computational graphs, we execute all computational graphs and filter out those that cannot be executed effectively, thereby obtaining answers for the valid computational graphs. Subsequently, we utilize LLMs to generate problems based on these graphs, as detailed in Appendix C. This method is similar to the approach human educators use when creating questions: they first consider the solution process and then formulate corresponding questions. Consequently, this approach enhances the grounding and logical coherence of the final synthesized problems.
@@ 3.4 Comparison to Previous Work
Table 2: After training on 50k data, the evaluation results of our method compared with various synthesis approaches and two datasets containing human-crafted problems across five benchmarks. The best and second-best ones among LLMs with the same backbone model are marked in bold and underlined respectively.
@@ 4.1 Experimental Setup
We compare our method with previous direct problem synthesis methods with publicly available datasets, including: (1) MetaMath Yu et al. (2024); (2) Orca-Math Mitra et al. (2024); (3) MathScale Tang et al. (2024);(4) Mammoth2 Yue et al. (2024); (5) Jiuzhang3.0 Zhou et al. (2024);(6) PromptCoT Zhao et al. (2025) and (7) ScaleQuest Ding et al. (2024). We also add two datasets containing human-crafted problems as strong baseline: (7) MMIQC Liu et al. (2024) contains QA pairs from Mathematics Stack Exchange and iteratively constructs more complex problems from MATH training data; (8) Numina-Math LI et al. (2024) is a large collection of both real-world human-crafted and synthesized math problems. We utilize the high-quality cn_k12, Olympiads, and synthetic_math subsets.
@@
Directly comparing the performance of models from previous works introduces potential unfair factors, as newer models may leverage more advanced models for backbone and answer generation. To ensure a fair comparison, we use the same model, Qwen-2.5-Math-7B-Instruct Yang et al. (2024) with temperature set to 0.7 and top-$p$ set to 0.95, to generate answers for problems synthesized by each method and train the same backbone model. By controlling for other variables, we can focus solely on comparing the quality of the synthesized problems.
@@
We assess the models’ performance on MATH-500 Hendrycks et al. (2021); Lightman et al. (2024), GSM8K Cobbe et al. (2021), GSM-Hard Gao et al. (2023), College Math Tang et al. (2024) and OlympiadBench He et al. (2024). The generated outputs are all in the form of natural language Chain-of-Thought (CoT) Wei et al. (2022) reasoning through greedy decoding, and we report zero-shot pass@1 accuracy.
@@
We extract seed data from the high-quality Numina-Math dataset LI et al. (2024) to collect function nodes and fine-tuning data. The problem decomposition is performed using Qwen2.5-72B-Instruct, resulting in a function library containing 50k nodes. We then use the fine-tuned Qwen2.5-7B-Instruct to generate computational graphs with the strategy described in 3.3.1, followed by problem back-translation using Qwen2.5-72B-Instruct.
@@ 4.2 Main Results
As shown in Table 2, methods focusing on synthesizing diverse new problems (e.g., Mammoth2, Jiuzhang3.0, ScaleQuest) outperform approaches that augment existing data (e.g., MetaMath, Orca-Math). This performance gap may be attributed to augmentation-based methods leading to new instructions that are too similar to the original ones, emphasizing the crucial role of data diversity. On the other hand, datasets containing real-world human-designed problems exhibit a clear performance advantage over existing synthetic data, highlighting the superior quality of human-designed problems. Our approach outperforms traditional data synthesis methods and yields results that closely match or even surpass the effectiveness of human-designed problems, without relying on proprietary models like GPT-4. Notably, this advantage is particularly evident on more challenging or computation-intensive datasets such as MATH-500, GSM-Hard, and OlympiadBench, further emphasizing the benefits of rational synthesis.
4.3 Further Analysis

### v2

@@ Abstract
The advancement of reasoning capabilities in Large Language Models (LLMs) requires substantial amounts of high-quality reasoning data, particularly in mathematics. Existing data synthesis methods, such as data augmentation from annotated training sets or direct question generation based on relevant knowledge points and documents, have expanded datasets but face challenges in mastering the internal logic of the problem during generation and ensuring the verifiability of the solutions. To address these issues, we propose RV-Syn, a novel Rational and Verifiable mathematical Synthesis approach. RV-Syn first constructs a structured library of mathematical operations and then composes them into executable computational graphs, which serve as verifiable solution blueprints. These graphs are subsequently back-translated into complex problems, enabling solution-guided, logic-aware problem generation while inherently ensuring the verifiability of the solving process. Experimental results show RV-Syn surpasses existing synthesis methods, including those involving human-crafted problems. Our method achieves a 6.3% performance gain over the previous state-of-the-art synthetic data on LLaMA-3-8B and demonstrates superior data efficiency, outperforming others with only half the training data (50k vs. 100k), enabling a more scalable and robust reasoning dataset generation framework.
@@
Jiapeng Wang1,3thanks: Equal contribution., Jinhao Jiang1,311footnotemark: 1, Zhiqiang Zhang2 , Jun Zhou2 , Wayne Xin Zhao1,3thanks: Corresponding author., 1Gaoling School of Artificial Intelligence, Renmin University of China 2Ant Group 3Beijing Key Laboratory of Research on Large Models and Intelligent Governance wangjp1010@ruc.edu.cn, batmanfly@gmail.com
@@ 1 Introduction
To address this scarcity, researchers have explored various synthesizing methods, particularly in the mathematics domain. The mainstream methods involve data augmentation based on existing annotated training sets, such as GSM8K Cobbe et al. (2021) and MATH Hendrycks et al. (2021), ranging from self-evolving instructions Xu et al. (2024a); Zeng et al. (2024) and question paraphrasing Yu et al. (2024), to solution augmentation Lu et al. (2024). However, these methods are limited by the available training data, constraining the synthesis diversity Li et al. (2024c, 2023). To enhance diversity, recent approaches enable LLMs to generate a large scale of questions from various mathematics-related sources, including web pages Yue et al. (2024) and knowledge points Tang et al. (2024) from web corpora or textbooks.
@@
However, as shown in Figure 1, these direct problem generation methods suffer from a fundamental limitation. We argue this stems from the auto-regressive nature of LLMs, which leads to a lack of forward-looking planning during generation. When composing a math problem, a model might introduce entities and quantities without a coherent plan, frequently resulting in internal contradictions, unsolvable scenarios, or trivial questions, particularly in complex, multi-step reasoning problems. Furthermore, this approach makes it difficult to validate the correctness of the generated solutions, often compromising the quality and reliability of the training data.
@@
Inspired by this, we propose RV-Syn, a novel framework for mathematical problem synthesis that emphasizes enhanced rationale and verifiability. Instead of direct generation, RV-Syn first synthesizes a structured computational graph. This graph serves as a formal blueprint, explicitly defining the problem’s reasoning structure and guaranteeing a valid, executable solution path. Subsequently, this logically consistent blueprint is back-translated into a natural language problem. By generating questions through solution guidance, RV-Syn effectively provides the rich semantics and control flows required for complex reasoning. As shown in Figure 1, this enables the synthesis of diverse and sophisticated reasoning structures—including multi-hop reasoning, iterative loops, and modular structures—while naturally providing verifiable ground-truth labels.
We conduct extensive experiments to evaluate the proposed RV-Syn method using various LLMs. The results demonstrate that our method achieves superior performance across five benchmarks compared to existing approaches, including those involving human-designed problems (e.g., Numina-Math). Notably, RV-Syn outperforms the previous state-of-the-art method while utilizing only half the data, leading to a more efficient scaling curve.
@@ 2 Related Work
Researchers have proposed various approaches to enhance the mathematical reasoning capabilities, including methods applied in training or inference stages. During the training stage, existing methods aim to enhance the LLMs from the aspects of pre-training and post-training. Specifically, some studies Azerbayev et al. (2024); Shao et al. (2024); Chen et al. (2024a) involve collecting extensive math-related corpora and enhancing the foundational mathematical capabilities of LLMs through continual pre-training. In contrast, other studies Chen et al. (2024c); Tong et al. (2024); Huang et al. (2024b); Toshniwal et al. (2024) focus on synthesizing a substantial amount of high-quality math-related instructional data, further refining the problem-solving abilities of LLMs through post-training. During the inference stage, there are two primary prompting approaches, including Chain-of-Thought (CoT) Wei et al. (2022) and Program-of-Thought (PoT) Chen et al. (2023). Specifically, some studies Yao et al. (2023); Shinn et al. (2023); Zhou et al. (2023) utilize CoT to elicit LLMs’ inherent reasoning ability with a step-by-step reasoning process. Furthermore, other studies Gou et al. (2024); Wang et al. (2024); Chen et al. (2024b); Wen et al. (2024) incorporate PoT to allow LLMs to utilize external computers during reasoning. In this work, we focus on the training stage and try to synthesize instruction data by leveraging the strengths of both the CoT and PoT methods.
@@
Existing research has proposed various synthesizing approaches: Firstly, existing studies focus on synthesizing new mathematical problems based on few-shot problems Li et al. (2024a), math-related documents Yue et al. (2024), key knowledge points Huang et al. (2024a), or self-evolution Luo et al. (2023). Furthermore, with the improvement of LLMs’ capabilities, subsequent studies Zhou et al. (2024); Ding et al. (2024) utilizingexplore the utilization of open-source LLMs (i.e., DeepSeek-Math and Qwen-Math) to scale the number of synthesized problems. Despite their effectiveness, direct problem generation suffers from a lack of fine-grained control over the problem’s internal logic, leading to lower-quality problems. Besides, these methods frequently lack mechanisms for validating the correctness of generated answers, which may introduce errors and affect the final performance. In contrast, our method starts from the internal logic of problem-solving, rather than directly synthesizing problems. This provides more refined logical consistency and improves problem quality, while also enabling effective correctness control.
@@ 3.1 Problem Decomposition
We first represent the solutions of existing mathematical problems with sequences of functions into a standardized format shown in Appendix C leveraging LLMs. Each function represents a specific mathematical skill with defined parameters, forming an executable computational graph that captures the data flow of the problem-solving process (examples shown in Figure 2). To ensure the correctness of the extracted solutions and their associated functions, we filter out erroneous cases by executing the solution code using a Python interpreter and comparing its output with the ground-truth answer provided in the corresponding Chain-of-Thought annotation. This results in a large-scale, high-quality set of computational graphs derived from real-world problems.
@@ 3.2 Graph-based Function Library Construction
After decomposing the problem, we extract functions from the computational graphs. To further enhance the quality of the function library, we employ model-based labeling. We utilize LLMs to validate the correctness of the functions and identify their corresponding mathematical topics, eliminating any functions that contain errors. The prompts are shown in Appendix A. To describe the relationships between these functions, each function is represented as a node, and they are organized into a graph-based function library. Specifically, the construction process involves two key steps: structural and semantic-aware function merging, followed by co-occurrence and topic-based node connecting.
3.2.1 Structural and Semantic-aware Function Merging
@@
Specifically, we examine the similarity between two functions by analyzing both structure and semantic aspects. For the structure, we focus on the computational operations without considering specific variable names. For instance, expressions such as “(a+b)/c(a+b)/c” and “(m+n)/d(m+n)/d” are considered equivalent. To accomplish this, we extract the function body and parse it using Abstract Syntax Trees (ASTs). For the semantic, we compute hash values for the function’s docstrings. If the ASTs or hash values of two functions are identical, we merge these functions into a single node. We show an example of two similar functions in Table 12 at Appendix H. This approach allows similar functions (e.g., matrix multiplication and matrix projection) to preserve their distinct mathematical meanings while minimizing redundancy in the final function library, offering advantages over direct deduplication. Our structure-based method using ASTs proves highly effective in practice, successfully identifying and merging approximately 28% of redundant functions from the initial library.
@@ 3.2.2 Co-occurrence and Topic-based Node Connecting
Specifically, for any two given nodes (i.e., two sets of functions), we first check whether there exists at least one pair of functions—one from each set—that co-occurred in the same computational graph. If so, we add a co-occurrence-based edge between these two nodes. If not, we then check whether there is at least one pair of functions sharing the same topic. If this condition is met, we add a topic-consistent edge. If neither condition is satisfied, no edge is added between the nodes.
In this way, we finally obtain an incompletely connected graph that contains multiple types of edges, which depicts the relationships between nodes. This serves as the basis for the subsequent recombination of functions.
@@ 3.3 Problem Synthesis
Based on the constructed graph-format function library, we are able to synthesize complex problems in a solution-guided approach. This process consists of two primary steps: first sampling various functions to regenerate the computational graph, then synthesizing the final complex problems.
@@ 3.3.1 Computational Graph Regeneration
To balance the reasonableness and novelty of the computational graphs, we design three sampling strategies: first, selecting nodes connected by co-occurrence-based edges; second, selecting nodes connected by topic-based edges; and third, selecting unconnected nodes. During the sampling process, nodes are randomly selected using these strategies. Since each node represents a set of functions, we further randomly select one function per node. Subsequently, we generate a computational graph based on these selected functions as detailed in Appendix A. This approach offers several advantages: Firstly, by combining functions, the generated computational graph can comprehensively cover mathematical skills while minimizing redundancy. Secondly, the sampling strategy effectively balances high-frequency combinations of mathematical skills (i.e., connected by edges) with long-tail combinations (i.e., without connected edges). In our setup, we randomly select from the three strategies when synthesizing each data sample.
3.3.2 Problem Back-translation and Verification
The final stage of our pipeline generates a complete (problem, solution) pair and ensures its correctness, highlighting the core verifiability of our framework. First, we execute all newly generated computational graphs using an interpreter. Graphs that fail to execute due to errors are discarded. This step serves two key purposes: it validates the feasibility of the synthesized reasoning process and yields a reliable, ground-truth final answer from successful execution. With a verified computational graph and its ground-truth answer, we then back-translate the graph into a coherent natural-language math problem and generate its Chain-of-Thought (CoT) solution. Unlike direct generation methods, which struggle to verify the correctness of their outputs, our framework can compare the final answer derived from the LLM-generated CoT solution against the ground-truth answer obtained from executing the computational graph. Any (problem, solution) pair with mismatched answers is automatically discarded. This execution-based filtering acts as a precise and automatic quality control mechanism, effectively eliminating instances of flawed reasoning or incorrect calculations in the final LLM-generated solutions. We provide more statistics of the synthesis pipeline in Appendix E.
@@ 3.4 Comparison to Previous Work
Table 2: After training on 50k data, the evaluation results of our method compared with various synthesis approaches across five benchmarks. The best ones among LLMs with the same backbone model are marked in bold.
@@ 4.1 Experimental Setup
We compare our method with previous problem synthesis methods with publicly available datasets, including: (1) MetaMath Yu et al. (2024) introduces several question bootstrapping techniques; (2) Orca-Math Mitra et al. (2024) augments existing datasets using an Agent-Instruct method; (3) MathScale Tang et al. (2024) uses topic and knowledge-point graphs to prompt new problem synthesis; (4) Mammoth2 Yue et al. (2024) extracts QA pairs from webpages; (5) Jiuzhang3.0 Zhou et al. (2024) uses math-related seed data to synthesize new QA pairs; (6) PromptCoT Zhao et al. (2025) generate complex problems with intermediate thought; (7) ScaleQuest Ding et al. (2024) trains a problem generation model similar to Magpie Xu et al. (2024b).
@@
Directly comparing the performance of models from previous works introduces potential unfair factors, as newer models may leverage more advanced models for backbone and answer generation (see Appendix  F for details). To ensure a fair comparison, we use the same model, Qwen-2.5-Math-7B-Instruct Yang et al. (2024) to generate answers for problems synthesized by each method and train the same backbone model. By controlling for other variables, we can focus on comparing the quality of the synthesized problems.
@@
We assess the models’ performance on MATH-500 Hendrycks et al. (2021); Lightman et al. (2024), GSM8K Cobbe et al. (2021), GSM-Hard Gao et al. (2023), College Math Tang et al. (2024) and OlympiadBench He et al. (2024). The generated outputs are all in the form of natural language Chain-of-Thought (CoT) Wei et al. (2022) through greedy decoding, and we report zero-shot pass@1 accuracy.
@@
We extract seed data from the high-quality Numina-Math dataset LI et al. (2024) to collect function nodes and fine-tuning data. The problem decomposition is performed using Qwen2.5-72B-Instruct. We then use the fine-tuned Qwen2.5-7B-Instruct to generate computational graphs with the strategy described in 3.3.1, followed by problem back-translation using Qwen2.5-72B-Instruct.
@@ 4.2 Main Results
As shown in Table 2, methods focusing on synthesizing diverse new problems (e.g., Mammoth2, Jiuzhang3.0, ScaleQuest) outperform approaches that augment existing data (e.g., MetaMath, Orca-Math). This performance gap may be attributed to augmentation-based methods leading to new instructions that are too similar to the original ones, emphasizing the crucial role of data diversity. Our approach outperforms traditional data synthesis methods without relying on proprietary models like GPT-4. For example, on LLaMA-3-8B-Instruct, our method outperforms the previous state-of-the-art by 6.3%. Notably, this advantage is particularly evident on more challenging or computation-intensive datasets such as MATH-500, GSM-Hard, and OlympiadBench, further emphasizing the benefits of rational synthesis. We present experiments on distilling Long-CoT reasoning in Appendix 4.7.
4.3 Scalability Study